An organic compound 0.246g by mass was burnt in oxygen and produced 0.198g carbon dioxide and 0.1014g water what is the empirical formula of this compound

1
Expert's answer
2022-03-13T15:19:47-0400

Solution:

Convert to moles:

The molar mass of CO2 is 44.01 g/mol

Therefore,

Moles of CO2 = (0.198 g CO2) × (1 mol CO2 / 44.01 g CO2) = 0.00450 mol CO2

Schematic equation:

CO2 → C

According to the shematic equation above:

Moles of C = Moles of CO2 = 0.00450 mol

Moles of C = 0.00450 mol

The molar mass of C is 12.0107 g/mol

Therefore,

Mass of C = (0.00450 mol C) × (12.0107 g C / 1 mol C) = 0.054 g C

 

The molar mass of H2O is 18.0153 g/mol

Therefore,

Moles of H2O = (0.1014 g H2O) × (1 mol H2O / 18.0153 g H2O) = 0.00563 mol H2O

Schematic equation:

H2O → 2H

According to the shematic equation above

Moles of H = 2 × Moles of H2O = 2 × 0.00563 mol = 0.01126 mol

Moles of H = 0.01126 mol

The molar mass of H is 1.00784 g/mol

Therefore,

Mass of H = (0.01126 mol H) × (1.00784 g H / 1 mol H) = 0.0113 g H


Mass of H + Mass of C = 0.0113 g + 0.054 g = 0.0653 g

(Mass of H + Mass of C) < (Mass of organic compound)

Thus, the compound contains another element besides oxygen and hydrogen.

Suppose the remaining element is bromine (Br).

Then,

Mass of Br = Mass of organic compound − (Mass of H + Mass of C)

Mass of Br = 0.246 g − 0.0653 g = 0.1807 g

The molar mass of Br is 79.904 g/mol

Therefore,

Moles of Br = (0.1807 g Br) × (1 mol Br / 79.904 g Br) = 0.00226 mol Br

Moles of Br = 0.00226 mol


Divide all moles by the smallest of the results:

C: 0.00450 mol / 0.00226 mol = 1.99 = 2

H: 0.01126 mol / 0.00226 mol = 4.98 = 5

Br: 0.00226 mol / 0.00226 mol = 1


Thus, the empirical formula of organic compound is C2H5Br


Answer: The empirical formula of the compound is C2H5Br